Please help me solve this equation: show me how to calculate
x^1+x^2+x^3=0
2x^1-x^2-x^3=-3
x^1-x^2+x^3=0
Answer= x^-1
x^0
x^1

Answer by Alan3354(69443)

x^1+x^2+x^3=0
2x^1-x^2-x^3=-3
x^1-x^2+x^3=0
-----------------
Eqns 1 and 3 --> x^2 = 0
If x^2 = 0, then all the terms are zero, and the 2nd eqn is not possible.
You sure those are exponents? Should they be subscripts?